# Maximal Inequalities for Separately Exchangeable Empirical Processes

### v1

@@ Abstract.
This paper derives new maximal inequalities for empirical processes associated with separately exchangeable random arrays. For any fixed index dimension $K\geq 1$, we establish a global maximal inequality that bounds the $q$-th moment, for any $q\in[1,\infty)$, of the supremum of these processes. In addition, we obtain a refined local maximal inequality that controls the first absolute moment of the supremum. Both results are proved for a general pointwise measurable class of functions.
@@ 1. Introduction
The local maximal inequality is a crucial tool for obtaining sharp convergence rates. Apart from the classical i.i.d. case ($K=1$), no local maximal inequality for SE arrays exists in the literature. Establishing a local maximal inequality for SE arrays is especially challenging because their intricate multiway dependence structure induces complex interactions among observations. Unlike in the i.i.d. or $U$-statistics settings, the multidimensional dependencies inherent to SE arrays render classical techniques—such as symmetrisation and Hoeffding averaging—inapplicable. To address these challenges, we propose a novel proof strategy that partitions the index set into transversal groups. This innovative construction effectively decouples the dependencies among observations within each group, thereby enabling the application of the Hoffmann–Jørgensen inequality to derive sharp bounds for terms involving higher moments. Consequently, our work fills a critical gap by providing both global and local maximal inequalities for a potentially uncountable but pointwise measurable class of functions under SE sampling.
@@ 2. Main Results
Finally, fix any $1\leq k\leq K$ and $\bm{e}\in\mathcal{E}_{k}$. Define the uniform entropy integral by
@@
The following result is a general global maximal inequality for SE empirical processes with an arbitrary index order $K$ and for a general order of moment $q\in[1,\infty)$. Its proof follows the arguments in the proof of Corollary B.1 in Chiang et al. (2023) with some modifications to account for a general class of functions.
@@
Suppose $\mathcal{F}:\mathcal{S}\to\mathbb{R}$ is a pointwise measurable class of functions. Let $(X_{\bm{i}})_{{\bm{i}}\in[\bm{N}]}$ be a sample from $S$-valued separately exchangeable random vectors $(X_{\bm{i}})_{{\bm{i}}\in\mathbb{N}^{K}}$. Pick any $1\leq k\leq K$ and $\bm{e}\in\mathcal{E}_{k}$. Then, for any $q\in[1,\infty)$, we have
@@
$\left(\mathbb{E}\left[\left\|\frac{1}{\sqrt{|I_{{\bm{N}},{\bm{e}}}|}}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}(\pi_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}^{q}\right]\right)^{1/q}$
@@
Although the global maximal inequality works for general $q$, in the case that the supremum of the first absolute moment is concerned, local maximal inequalities usually provides shaper bounds. The following is a novel local maximal inequality for SE empirical processes. Unlike the proof of the global maximal inequality, which is largely analogous to the corresponding results for $U$-processes that can be found in Chen and Kato (2019), its proof relies on a novel argument that utilizes construction of a partition with the property that each block in the partition satisfies a transversality property. We present this construction in Lemma 1 below.
@@
Suppose $\mathcal{F}:\mathcal{S}\to\mathbb{R}$ is a pointwise measurable class of functions. Let $(X_{\bm{i}})_{{\bm{i}}\in[\bm{N}]}$ be a sample from $S$-valued separately exchangeable random vectors $(X_{\bm{i}})_{{\bm{i}}\in\mathbb{N}^{K}}$. Set ${\bm{e}}\in\{0,1\}^{K}$ and let $\sigma_{\bm{e}}$ be a constant such that $\sup_{f\in\mathcal{F}}\|P_{\bm{e}}f\|_{P,2}\leq\sigma_{\bm{e}}\leq\|P_{\bm{e}}F\|_{P,2}$,
@@
$\mathbb{E}[\|H_{\bm{N}}^{\bm{e}}(f)\|_{\mathcal{F}}]\lesssim$ $\mathbb{E}\left[\left|\frac{1}{|I_{{\bm{N}},{\bm{e}}}|}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}\varepsilon_{1,i_{1}}...\varepsilon_{k,i_{k}}(\pi_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}\right].$
@@
$\mathbb{E}\left[\left|\frac{1}{|I_{{\bm{N}},{\bm{e}}}|}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}\varepsilon_{1,i_{1}}...\varepsilon_{k,i_{k}}(P_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}\right]$
@@
Thus $z\leq\widetilde{z}:=\sqrt{B_{n,{\bm{e}}}}/\|(\pi_{\bm{e}})F\|_{P,2}$. Note that by symmetrisation inequality for independent processes, the contration principle (Theorem 4.12. in Ledoux and Talagrand 1991), and the Cauchy-Schwartz inequality, one has
@@
Our goal is to partition $[{\bm{N}}]$ into subsets (which we call groups) of size $N_{K}$ that are transversal. For each $j=1,2,\dots,K-1$, define the function $\alpha_{j}\colon[N_{j}]\to\{0,1,2,\dots\}\quad\text{by}\quad\alpha_{j}(g)=g-1.$ Then, for $j=1,2,\dots,K-1$, define $\phi_{j}\colon[N_{K}]\times[N_{j}]\to[N_{j}]$ by
$\phi_{j}(t,g)=\Bigl{(}((t+g-2)\mod N_{j})+1\Bigr{)}.$
@@
Thus, each group contains $N_{K}$ elements. We now claim that groups based on this mapping forms a partition and satisfies transversality in each group.

### v2

@@ Abstract.
This paper derives new maximal inequalities for empirical processes associated with separately exchangeable random arrays. For fixed index dimension $K\geq 1$, we establish a global maximal inequality bounding the $q$-th moment ($q\in[1,\infty)$) of the supremum of these processes. We also obtain a refined local maximal inequality controlling the first absolute moment of the supremum. Both results are proved for a general pointwise measurable function class. Our approach uses a new technique partitioning the index set into transversal groups, decoupling dependencies and enabling more sophisticated higher moment bounds.
@@ 1. Introduction
The local maximal inequality is a crucial tool for obtaining sharp convergence rates. Apart from the classical i.i.d. case ($K=1$), no local maximal inequality for SE arrays exists in the literature. Establishing a local maximal inequality for SE arrays is especially challenging because their intricate multiway dependence structure induces complex interactions among observations. Unlike in the i.i.d. or $U$-statistics settings, the multidimensional dependencies inherent to SE arrays render classical techniques—such as symmetrisation and Hoeffding averaging—inapplicable. To address these challenges, we propose a novel proof strategy that carefully partitions the index set into transversal groups. This innovative construction effectively decouples the dependencies among observations within each group, thereby enabling the application of the Hoffmann–Jørgensen inequality to derive sharp bounds for terms involving higher moments. Consequently, our work fills a critical gap by providing both global and local maximal inequalities for a potentially uncountable but pointwise measurable class of functions under SE sampling.
@@ 2. Main Results
Finally, fix any $1\leq k\leq K$ and $\bm{e}\in\mathcal{E}_{k}$. For a given class $\mathcal{F}$ with an envelope $F$, define for a $\delta>0$ its uniform entropy integral by
@@
The following result is a general global maximal inequality for SE empirical processes with an arbitrary index order $K$ and for a general order of moment $q\in[1,\infty)$. Its proof follows the arguments in the proof of Corollary B.1 in Chiang et al. (2023) with some modifications to account for a more general class of functions.
@@
Suppose $\mathcal{F}:\mathcal{S}\to\mathbb{R}$ is a pointwise measurable class of functions with an envelope $F$. Let $(X_{\bm{i}})_{{\bm{i}}\in[\bm{N}]}$ be a sample from $S$-valued separately exchangeable random vectors $(X_{\bm{i}})_{{\bm{i}}\in\mathbb{N}^{K}}$. Pick any $1\leq k\leq K$ and $\bm{e}\in\mathcal{E}_{k}$. Then, for any $q\in[1,\infty)$, we have
@@
$|I_{{\bm{N}},{\bm{e}}}|^{1/2}\left(\mathbb{E}[\|H_{\bm{N}}^{\bm{e}}(f)\|_{\mathcal{F}}^{q}]\right)^{1/q}=$ $\left(\mathbb{E}\left[\left\|\frac{1}{\sqrt{|I_{{\bm{N}},{\bm{e}}}|}}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}(\pi_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}^{q}\right]\right)^{1/q}$
@@
Although the global maximal inequality works for general $q$, in the case that the supremum of the first absolute moment is concerned, local maximal inequalities usually provides shaper bounds. The following is a novel local maximal inequality for SE empirical processes. Unlike the proof of the global maximal inequality, which is largely analogous to the corresponding results for $U$-processes that can be found in Chen and Kato (2019), its proof relies on a novel argument that utilises construction of a partition with the property that each block in the partition satisfies a transversality property. We present this construction in Lemma 1 below.
@@
Suppose $\mathcal{F}:\mathcal{S}\to\mathbb{R}$ is a pointwise measurable class of functions with an envelope $F$. Let $(X_{\bm{i}})_{{\bm{i}}\in[\bm{N}]}$ be a sample from $S$-valued separately exchangeable random vectors $(X_{\bm{i}})_{{\bm{i}}\in\mathbb{N}^{K}}$. Set ${\bm{e}}\in\{0,1\}^{K}$ and let $\sigma_{\bm{e}}$ be a constant such that $\sup_{f\in\mathcal{F}}\|P_{\bm{e}}f\|_{P,2}\leq\sigma_{\bm{e}}\leq\|P_{\bm{e}}F\|_{P,2}$,
@@
$|I_{{\bm{N}},{\bm{e}}}|^{1/2}\mathbb{E}[\|H_{\bm{N}}^{\bm{e}}(f)\|_{\mathcal{F}}]\lesssim$ $\mathbb{E}\left[\left|\frac{1}{\sqrt{|I_{{\bm{N}},{\bm{e}}}|}}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}\varepsilon_{1,i_{1}}...\varepsilon_{k,i_{k}}(\pi_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}\right].$
@@
$\mathbb{E}\left[\left|\frac{1}{\sqrt{|I_{{\bm{N}},{\bm{e}}}|}}\sum_{{\bm{i}}\in I_{{\bm{N}},{\bm{e}}}}\varepsilon_{1,i_{1}}...\varepsilon_{k,i_{k}}(P_{{\bm{e}}}f)(\{U_{{\bm{i}}\odot{\bm{e}}^{\prime}}\}_{{\bm{e}}^{\prime}\leq{\bm{e}}})\right\|_{\mathcal{F}}\right]$
@@
Thus $z\leq\widetilde{z}:=\sqrt{B_{n,{\bm{e}}}}/\|(\pi_{\bm{e}})F\|_{P,2}$. Note that by symmetrisation inequality for independent processes, the contration principle (Theorem 4.12. in Ledoux and Talagrand 1991), and the Cauchy-Schwarz inequality, one has
@@
Our goal is to partition $[{\bm{N}}]$ into subsets (which we call groups) of size $N_{K}$ that are transversal. For $j=1,2,\dots,K-1$, define $\phi_{j}\colon[N_{K}]\times[N_{j}]\to[N_{j}]$ by
$\phi_{j}(t,g)=((t+g-2)\mod N_{j})+1.$
@@
Thus, each group contains $N_{K}$ elements.